OpenStudy (anonymous):

For x ≠ 0, x ≠ 1, and x ≠ −1 which of the expressions below is equivalent to the following. 1/x/1/x^2-1 choices are ~ 1/x-x^3 ~x/1-x^2 ~1/1-x^2 ~x/1-x ~1/x^3-x

OpenStudy (helder_edwin):

\[ \LARGE \frac{\frac{1/x}{1}}{x^2-1} \]

OpenStudy (anonymous):

thats not even one of the choices.

OpenStudy (helder_edwin):

i am rewritting what u have in order for u to do the operations correctly!!!!!! or u r just waiting for the answer??

OpenStudy (anonymous):

Oh okay. You can show me how to do it.

OpenStudy (helder_edwin):

first \[ \large \frac{a}{1}=? \]

OpenStudy (anonymous):

a

OpenStudy (helder_edwin):

yes so in the numerator what do u get??

OpenStudy (anonymous):

a

OpenStudy (helder_edwin):

no in the problem u r supposed to solve

OpenStudy (helder_edwin):

my very first post

OpenStudy (anonymous):

oh sorry I would get x

OpenStudy (helder_edwin):

no

OpenStudy (anonymous):

thats what I'm screwing up on. I don't know if I'm suppose to do the reciprocal and multiply that by the numerator.

OpenStudy (anonymous):

?

OpenStudy (helder_edwin):

\[ \large \frac{1/x}{1}=? \]

OpenStudy (helder_edwin):

\[ \large \frac{a}{1}= \] it is the same!!!

OpenStudy (anonymous):

thats not even the equation.

OpenStudy (helder_edwin):

u r not solving an equation. u r simplifying an expresion

OpenStudy (anonymous):